# NuSTAR discovers a long type-I X-ray burst from the clocked burster GS 1826-24

### v1

@@ Abstract
The source GS 1826-24 is a neutron star low mass X-ray binary known as the “clocked burster” because of its extremely regular bursting behavior. We report on the detection of a long type-I X-ray burst from this source. We perform a detailed spectroscopic analysis of the long X-ray burst, lasting for s, seen in the NuSTAR observation carried out on 2022 September. The persistent emission is well described by an absorbed thermal Comptonization model nthcomp, and the source exhibits a hard spectral state during this observation. The observed burst exhibits a rise time of s and a decay time of s. The time-resolved spectroscopy of the burst shows a significant departure from a pure thermal spectrum and is described with a model consisting of a varying-temperature blackbody plus an evolving persistent emission component. We observe a significant enhancement in the persistent emission during the burst. The enhancement of the pre-burst persistent flux is possibly due to Poynting-Robertson drag or coronal reprocessing. At the peak of the burst, the blackbody temperature and the blackbody emitting radius reached a maximum of keV and km, respectively. The peak flux () during the burst is ergs cm-2 s-1, which corresponds to a luminosity of ergs s-1.
@@ 1 Introduction
The source was not active in its clocked burster state, and the pattern of regular bursting behaviour has been broken, with the transition to weaker, irregular bursts accompanied by a soft spectral state, in 2014 June (Chenevez et al., 2016). Shortly after this soft episode, the source returned to its hard state again. According to observations by Swift/BAT and MAXI, the intensity of GS 1826-238 gradually decreased and transitioned back to the hard state (Chenevez et al., 2016, Ji et al., 2018). After 2016, this source transfers into intermediate and soft states (Ji et al., 2018). There have been 15 NuSTAR StrayCat observations of GS 1826-24 between 2014 February and 2021 November, and a transition of the source from the island atoll state (hard) to a banana branch (soft) has been observed during this time interval (Yun et al., 2023). Again, approximately after one year, NuSTAR observed the source in 2022 September in a typical hard spectral state, and here we present the analysis of this NuSTAR observation. Presumably, the change in the observed behavior is correlated to a change in the mass accretion rate onto the NS (Hasinger and van der Klis, 1989). Recent observation suggests that GS 1826-238 may have entered again a clocked burster state for the first time in 10 years with a shortest burst recurrence period of approximately 1.6 hours (Iwata et al., 2025).
Chenevez et al. (2016) detected the Eddington limited type-I X-ray burst and derived the distance of the source to be kpc, assuming an isotropic emission. An excess in the soft and hard X-ray flux was reported during the type-I X-ray burst observed in this source (Ji et al. 2014, Sánchez-Fernández et al. 2020). The source is also known for the long burst tails, lasting s, which were powered by rp-process hydrogen burning (Galloway et al., 2004). The inferred accretion rate varies little on short timescales, and is within the range (Galloway et al. 2008, Chenevez et al. 2016). The NICER observations of this source revealed 7–9 mHz oscillations with fractional rms 2 per cent at 6 keV (Strohmayer et al., 2018). The first superburst (lasted for at least 3 hours) from GS 1826–24 was detected by the MAXI/GSC on 2018 February (Iwakiri et al., 2018).
@@ 3 Light Curve
In the left panel of Figure 1, we show the NuSTAR/FPMA light curve of the source in the keV energy band. The average count rate during the NuSTAR observation was counts s-1. One type-I thermonuclear X-ray burst is present at about th ks (MJD ) during the NuSTAR observation (see left and middle panel of Figure 1). From this particular NuSTAR orbit, a change in the persistent flux has been observed (see right panel of Figure 1). The X-ray burst has a typical sharp rise and near-exponential decay. The observed burst has a long tail, lasting s, and the peak of the burst seems to reach approximately counts s-1, about a factor of above the level of the persistent emission. Considering the rising and exponential decay time, the type-I burst lasts about s. We extracted the light curve (see Figure 2) of the burst in different energy bands. We performed the time-resolved spectral analysis of the burst to investigate the spectral properties in detail. For the persistent emission spectrum, we took a time interval of around s just before the rise of the burst and a time interval of 1 ks before the burst, but from a different NuSTAR orbit.
@@ 4.2 Burst spectral analysis
To take into account the residuals of the blackbody burst fit, we employed the enhanced persistent emission model. Initially, we performed the fit assuming that the persistent emission spectrum does not change during the burst, corresponding to the standard burst fitting technique (e.g. Kuulkers et al. 2002, Galloway et al. 2008). This model failed to fit the observed hard X-ray deficit and the soft X-ray excess for all the spectra. Later, we assumed that the amplitude of the persistent emission can change during the burst (Worpel et al., 2013) and adopted the following best-fit model composed of TBabs*(bbodyrad+ nthcomp). The blackbody and nthcomp models account for the X-ray burst and persistent emissions during the burst, respectively. The parameters of the nthcomp were fixed to the best-fit values of the preburst persistent emission (reported in Table 1). The parameter is a free scaling factor that can be used to account for the variation of the persistent emission. If , this means that the amplitude of persistent emission during the burst is equal to the value before the burst. This model improved the fit and provided a reasonable description of all the burst spectra (see bottom panels of Figure 4). We obtained an acceptable fit for all the intervals of the burst with the model TBabs*(bbodyrad+nthcomp), and it constrained all the free parameters simultaneously. We found that the best-fit values of is greater than for all segments, suggesting that the persistent emission typically increases during the burst. The best-fit parameter values for individual segments are quoted in Table 2. We used the XSPEC model cflux to determine the unabsorbed fluxes of different spectral components of the time-resolved spectra in the energy range keV during the burst (see Table 2).
@@ 5 Discussion
We performed a spectroscopic analysis of the long X-ray burst, observable for s, seen in the NuSTAR observation of the source GS 1826-24 to investigate the properties of the burst in detail. We first considered the spectral studies of the persistent emission of the source. We extracted persistent spectrum in the keV range from a time interval of 400 s just before the onset of the burst and a time interval of 2 ks before the burst, but from a different NuSTAR orbit. We modeled the persistent spectra using an absorbed cutoff power law (cutoffpl) or a thermal Comptonization model (nthcomp). Both models yielded a reasonably good description of the persistent emission in both cases. During this observation keV persistent flux was ergs cm-2 s-1, which correspond to a luminosity of ergs s-1 (assuming a source distance of kpc). This luminosity is roughly of the Eddington luminosity (where ergs s-1 for a NS). The inferred persistent flux is consistent with the hard state bolometric flux of the source (Thompson et al., 2008, Sánchez-Fernández et al., 2020). The local accretion rate per unit area can be calculated from the persistent luminosity following Galloway et al. (2008), which is g cm-2 s-1 for a NS of mass and radius 10 km in our case. Thus, the accretion rate onto the source was modest and exhibited a hard spectral state during this observation.
@@
We extracted seven separate spectra along the burst profile, in the intervals 20 s, 20 s, 30 s, 40 s, 50 s, 100 s, and 200 s from the burst onset. We initially modeled the burst spectra using a simple absorbed blackbody model, but excesses are visible at low and high energies. To fit the time-resolved spectra of the burst properly, we used a model consisting of a blackbody plus a variable persistent emission component. We accounted for the excess flux in X-ray burst emission using the model of Worpel et al. (2013). Along the burst profile, the blackbody temperature decreases from the peak value keV to keV during the tail of the burst. The blackbody emitting radius () is highest km at the peak of the burst and becomes smaller as the flux decreases, but not in a regular fashion (see Figure 5). We did not observe any indication of photospheric radius expansion (PRE) at the peak of the burst from the time-resolved spectroscopy. The peak flux () during the burst is ergs cm-2 s-1. This flux corresponds to a luminosity of ergs s-1, which is roughly of the (consistent with the typical hard state bursts), indicating that the flux at the peak of the burst did not reach the Eddington limit. Additionally, we found that the best-fit values of are greater than , suggesting that the persistent emission typically increases during the burst. The values rise at the burst start and quickly reach the maximum () within seconds, then they decrease slowly and return to around unity during the cooling tail (see top panel of Figure 5). This indicates that the persistent emission is rapidly enhanced due to the X-ray burst radiation and then returns to the preburst level during the cooling tail. The enhancement in is perfectly consistent with the results obtained by Ji et al. (2015) after analyzing burst spectra seen by RXTE. We determined the unabsorbed keV blackbody flux (), the persistent flux (), and showed their variation during the burst in the Figure 6. Around the burst peak, the blackbody flux supersedes the persistent flux; later on, the persistent flux starts dominating.
@@
The total duration of the burst, i.e., the time to evolve away from and return to the persistent state, was s in the NuSTAR FPMA/FPMB lightcurves. It is consistent with the fact that the hard state bursts have a longer burst duration and a larger burst fluence compared to the soft state bursts (Yun et al., 2023). During this burst duration, the intensity of the source increased by at least (at the beginning and end point of the burst) compared to its persistent intensity level. We found that at about seconds after the burst onset, the flux reached of its peak flux. The burst exhibited a moderate rise time of s, which is the time spent between the start of the burst and the point at which of the peak burst intensity is reached. The bolometric flux ( keV) during the burst is ergs cm-2 s-1 which has been used to calculate the burst fluence (). The value of was found to be ergs cm-2. The corresponding decay time of the burst is s. The measured fluence of the burst corresponds to a net burst energy release of erg (assuming a distance of kpc). The ignition depth () is related to the energy release during the burst by the equation , where is the energy release during nuclear reaction (Cumming, 2003). To calculate , we take an energy release of burning He to Ni of erg g-1 (including losses due to neutrino emission), and assume a NS mass of and radius of km, which give a gravitational redshift of . We found an ignitation depth g cm-2. At an accretion rate of g cm-2 s-1, the recurrence time corresponding to a column depth of g cm-2 is days. However, this can not be compared to the earlier recurrence times as the current burst is observed when the source was not in an active clocked burster state. We further calculated the factor to understand the burst fuel composition, which is the ratio of the persistent fluence between bursts to the total burst fluence of the burst, . We found , which exceeds nominal values for pure helium bursts by a factor (Galloway et al., 2008). However, in our case, we note that the is not less than one day.
@@
The burst fuel composition can be verified in another way by determining the local accretion rate (Galloway et al., 2008). A convenient unit of accretion rate is the Eddington rate. We, therefore, expressed the local accretion rate () as a fraction of the local Eddington accretion rate (). We assumed the local accretion rate g cm-2 s-1, then we obtained as a fraction of for (for solar abundance) and (for pure helium). We found that and for and , respectively. The value of less than for supports the idea that these bursts were ignited in a helium-rich environment (see e.g., Galloway and Keek 2021) and such accretion rate is expected based on the burst behavior (Galloway et al., 2008). Despite the uncertainty in the fuel composition, the properties of the bursts (i.e., longer burst duration) from GS 1826-24 strongly suggest burning of mixed H/He fuel by unstable He ignition (Fujimoto et al., 1981).
@@ 7 Acknowledgements
This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Centre (HEASARC). This research also has made use of the NuSTAR data analysis software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). ASM would like to thank Inter-University Centre for Astronomy and Astrophysics (IUCAA) for their facilities extended to him under their Visiting Associate Programme.

### v2

@@ Abstract
The source GS 1826-24 is a neutron star low mass X-ray binary known as the “clocked burster” because of its extremely regular bursting behavior. We report on the detection of a long type-I X-ray burst from this source. We perform a detailed spectroscopic analysis of the long X-ray burst, lasting for s, seen in the NuSTAR observation carried out on 2022 September. The persistent emission is well described by an absorbed thermal Comptonization model nthcomp, and the source exhibits a soft spectral state during this observation. The observed burst exhibits a rise time of s and a decay time of s. The time-resolved spectroscopy of the burst shows a significant departure from a pure thermal spectrum and is described with a model consisting of a varying-temperature blackbody plus an evolving persistent emission component. We observe a significant enhancement in the persistent emission during the burst. The enhancement of the pre-burst persistent flux is possibly due to Poynting-Robertson drag or coronal reprocessing. At the peak of the burst, the blackbody temperature and the blackbody emitting radius reached a maximum of keV and km, respectively. The peak flux () during the burst is ergs cm-2 s-1, which corresponds to a luminosity of ergs s-1.
@@ 1 Introduction
The source was not active in its clocked burster state, and the pattern of regular bursting behaviour has been broken, with the transition to weaker, irregular bursts accompanied by a soft spectral state, in 2014 June (Chenevez et al., 2016). Shortly after this soft episode, the source returned to its hard state again. According to observations by Swift/BAT and MAXI, the intensity of GS 1826-238 gradually decreased and transitioned back to the hard state (Chenevez et al., 2016, Ji et al., 2018). After 2016, this source transfers into intermediate and soft states (Ji et al., 2018). There have been 15 NuSTAR StrayCat observations of GS 1826-24 between 2014 February and 2021 November, and a transition of the source from the island atoll state (hard) to a banana branch (soft) has been observed during this time interval (Yun et al., 2023). Again, approximately after one year, NuSTAR observed the source in 2022 September in an extended soft spectral state, and here we present the analysis of this NuSTAR observation. Presumably, the change in the observed behavior is correlated to a change in the mass accretion rate onto the NS (Hasinger and van der Klis, 1989). Recent observation suggests that GS 1826-238 may have entered again a clocked burster state for the first time in 10 years with a shortest burst recurrence period of approximately 1.6 hours (Iwata et al., 2025).
Chenevez et al. (2016) detected the Eddington limited type-I X-ray burst and derived the distance of the source to be kpc, assuming an isotropic emission. An excess in the soft flux and a shortage in the hard X-ray flux was reported during the type-I X-ray burst observed in this source (Ji et al. 2014, Sánchez-Fernández et al. 2020). The source is also known for the long burst tails, lasting s, which were powered by rp-process hydrogen burning (Galloway et al., 2004). The inferred accretion rate varies little on short timescales, and is within the range (Galloway et al. 2008, Chenevez et al. 2016). The NICER observations of this source revealed 7–9 mHz oscillations with fractional rms 2 per cent at 6 keV (Strohmayer et al., 2018). The first superburst (lasted for at least 3 hours) from GS 1826–24 was detected by the MAXI/GSC on 2018 February (Iwakiri et al., 2018).
@@ 3 Light Curve
In the left panel of Figure 1, we show the NuSTAR/FPMA light curve of the source in the keV energy band. The average count rate during the NuSTAR observation was counts s-1. One type-I thermonuclear X-ray burst is present at about th ks (MJD ) during the NuSTAR observation (see left and middle panel of Figure 1). From this particular NuSTAR orbit, a change in the persistent flux has been observed (see right panel of Figure 1). The X-ray burst has a typical sharp rise and near-exponential decay. The observed burst has a long tail, lasting s, and the peak of the burst seems to reach approximately counts s-1, about a factor of above the level of the persistent emission. Considering the rising and exponential decay time, the type-I burst lasts about s. We extracted the light curve (see Figure 2) of the burst in different energy bands. We performed the time-resolved spectral analysis of the burst to investigate the spectral properties in detail. For the persistent emission spectrum, we took a time interval of around s just before the rise of the burst and a time interval of ks before the burst, but from a different NuSTAR orbit.
@@ 4.2 Burst spectral analysis
To take into account the residuals of the blackbody burst fit, we employed the enhanced persistent emission model. Initially, we performed the fit assuming that the persistent emission spectrum does not change during the burst, corresponding to the standard burst fitting technique (e.g. Kuulkers et al. 2002, Galloway et al. 2008). This model failed to fit the observed hard X-ray deficit and the soft X-ray excess for all the spectra. Later, we assumed that the amplitude of the persistent emission can change during the burst (Worpel et al., 2013) and adopted the following best-fit model composed of TBabs*(bbodyrad+ nthcomp). The blackbody and nthcomp models account for the X-ray burst and persistent emissions during the burst, respectively. The parameters of the nthcomp were fixed to the best-fit values of the preburst persistent emission taken from s time interval (reported in Table 1). The parameter is a free scaling factor that can be used to account for the variation of the persistent emission. If , this means that the amplitude of persistent emission during the burst is equal to the value before the burst. This model improved the fit and provided a reasonable description of all the burst spectra (see bottom panels of Figure 4). We obtained an acceptable fit for all the intervals of the burst with the model TBabs*(bbodyrad+nthcomp), and it constrained all the free parameters simultaneously. We found that the best-fit values of is greater than for all segments, suggesting that the persistent emission typically increases during the burst. The best-fit parameter values for individual segments are quoted in Table 2. We used the XSPEC model cflux to determine the unabsorbed fluxes of different spectral components of the time-resolved spectra in the energy range keV during the burst (see Table 2).
@@
We tested other scenarios to fit the burst spectra apart from the model. First, we attempted to fit the burst spectra by the model TBabs*thcomp*bbodyrad, considering that the burst photons could be affected by the corona/boundary layer (Chen et al., 2022). Here, thcomp (Zdziarski et al., 2020) is a more accurate version of nthcomp, and the parameter values of this model are extracted from the persistent fit. However, this model failed to fit the residuals observed at low and high energies and provided a fit to the burst spectra that was significantly worse compared to the previous model. Second, we tried the model TBabs*(bbodyrad+powerlaw+nthcomp) to fit the burst spectrum, in which the nthcomp parameter values were fixed to the persistent emission (Jaisawal et al., 2025). However, the current data failed to constrain all the free parameters simultaneously and could not produce conclusive results. Third, to fit the burst spectra, another combination of models like TBabs*(bbodyrad+
relxillNS+nthcomp) can be tested in which relxillNS accounts for the reflection of the burst photons from the accretion disk (Lu et al., 2023). However, we could not implement this model because of the short exposure time of each segment of the burst spectrum and the lack of a priori knowledge of some important disk-related parameters. The above-mentioned facts restricted us from discussing these models.
@@ 5 Discussion
We performed a spectroscopic analysis of the long X-ray burst, observable for s, seen in the NuSTAR observation of the source GS 1826-24 to investigate the properties of the burst in detail. We first considered the spectral studies of the persistent emission of the source. We extracted persistent spectrum in the keV range from a time interval of 400 s just before the onset of the burst and a time interval of 2 ks before the burst, but from a different NuSTAR orbit. We modeled the persistent spectra using an absorbed cutoff power law (cutoffpl) or a thermal Comptonization model (nthcomp). Both models yielded a reasonably good description of the persistent emission in both cases. During this observation keV persistent flux was ergs cm-2 s-1, which correspond to a luminosity of ergs s-1 (assuming a source distance of kpc). This luminosity is roughly of the Eddington luminosity (where ergs s-1 for a NS). The inferred persistent flux is consistent with the intermediate and soft state bolometric flux of the source (Ji et al., 2018, Sánchez-Fernández et al., 2020, Grefenstette et al., 2025). The local accretion rate per unit area can be calculated from the persistent luminosity following Galloway et al. (2008), which is g cm-2 s-1 for a NS of mass and radius 10 km in our case. Thus, the accretion rate onto the source was modest and exhibited a soft spectral state during this observation.
@@
We extracted seven separate spectra along the burst profile, in the intervals 20 s, 20 s, 30 s, 40 s, 50 s, 100 s, and 200 s from the burst onset. We initially modeled the burst spectra using a simple absorbed blackbody model, but excesses are visible at low and high energies. To fit the time-resolved spectra of the burst properly, we used a model consisting of a blackbody plus a variable persistent emission component. We accounted for the excess flux in X-ray burst emission using the model of Worpel et al. (2013). Along the burst profile, the blackbody temperature decreases from the peak value keV to keV during the tail of the burst. The blackbody emitting radius () is highest km at the peak of the burst and becomes smaller as the flux decreases, but not in a regular fashion (see Figure 5). We did not observe any indication of photospheric radius expansion (PRE) at the peak of the burst from the time-resolved spectroscopy. The peak flux () during the burst is ergs cm-2 s-1. This flux corresponds to a luminosity of ergs s-1, which is roughly of the , indicating that the flux at the peak of the burst did not reach the Eddington limit. Additionally, we found that the best-fit values of are greater than , suggesting that the persistent emission typically increases during the burst. The values rise at the burst start and quickly reach the maximum () within seconds, then they decrease slowly and return to around unity during the cooling tail (see top panel of Figure 5). This indicates that the persistent emission is rapidly enhanced due to the X-ray burst radiation and then returns to the preburst level during the cooling tail. To verify this, we further extracted a spectrum from a segment (after S7) of time interval s ( s) and fit in the same way as earlier for the spectra of segments S1…S7. The fit returned value of , thus confirming the above-mentioned fact. The enhancement in is perfectly consistent with the results obtained by Ji et al. (2015) after analyzing burst spectra seen by RXTE. We determined the unabsorbed keV blackbody flux (), the persistent flux (), and showed their variation during the burst in the Figure 6. Around the burst peak, the blackbody flux supersedes the persistent flux; later on, the persistent flux starts dominating.
@@
The total duration of the burst, i.e., the time to evolve away from and return to the persistent state, was s in the NuSTAR FPMA/FPMB lightcurves. The burst has characteristics of both short-duration soft-state bursts and long-duration hard-state bursts (Yun et al., 2023, Grefenstette et al., 2025). During this burst duration, the intensity of the source increased by at least (at the beginning and end point of the burst) compared to its persistent intensity level. We found that at about seconds after the burst onset, the flux reached of its peak flux. The burst exhibited a moderate rise time of s, which is the time spent between the start of the burst and the point at which of the peak burst intensity is reached. The bolometric flux ( keV) during the burst is ergs cm-2 s-1 which has been used to calculate the burst fluence (). The value of was found to be ergs cm-2. The corresponding decay time of the burst is s. The measured fluence of the burst corresponds to a net burst energy release of erg (assuming a distance of kpc). The ignition depth () is related to the energy release during the burst by the equation , where is the energy release during nuclear reaction (Cumming, 2003). To calculate , we take an energy release of burning He to Ni of erg g-1 (including losses due to neutrino emission), and assume a NS mass of and radius of km, which give a gravitational redshift of . We found an ignitation depth g cm-2. At an accretion rate of g cm-2 s-1, the recurrence time corresponding to a column depth of g cm-2 is days. However, this can not be compared to the earlier recurrence times as the current burst is observed when the source was not in an active clocked burster state.
@@
To understand the burst fuel composition, one can estimate the factor, which is the ratio of the persistent fluence between bursts to the total burst fluence of the burst, using the formula . In our case, there is no scope to calculate from the observation, as we see only a single burst in this observation. Therefore, we verified the burst fuel composition in another way by determining the local accretion rate (Galloway et al., 2008). A convenient unit of accretion rate is the Eddington rate. We, therefore, expressed the local accretion rate () as a fraction of the local Eddington accretion rate (). We assumed the local accretion rate g cm-2 s-1, then we obtained as a fraction of for (for solar abundance) and (for pure helium). We found that and for and , respectively. The value of less than for supports the idea that these bursts were ignited in a helium-rich environment (see e.g., Galloway and Keek 2021) and such accretion rate is expected based on the burst behavior (Galloway et al., 2008). Despite the uncertainty in the fuel composition, the properties of the bursts (i.e., longer burst duration) from GS 1826-24 strongly suggest burning of mixed H/He fuel by unstable He ignition (Fujimoto et al., 1981). However, to study various burst properties in great detail, we need an uninterrupted observation of this source with broadband X-ray coverage.
@@ 7 Acknowledgements
We want to thank the anonymous referee for his/her feedback and suggestions, which have significantly enhanced the quality of this paper. We would also like to thank Dr. Biplab Raychaudhuri for the productive discussions. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Centre (HEASARC). This research also has made use of the NuSTAR data analysis software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). ASM would like to thank Inter-University Centre for Astronomy and Astrophysics (IUCAA) for their facilities extended to him under their Visiting Associate Programme.